OpenStudy (nincompoop):

what condition would satisfy 2+2=5

OpenStudy (anonymous):

if mathmatics was faulty

OpenStudy (anonymous):

is this a serious question?

OpenStudy (anonymous):

reported.

OpenStudy (nincompoop):

@HawkCrimson this is a serious question

OpenStudy (lgbasallote):

this is possible if you're a mathematician. If you're an engineer, it's not possible

OpenStudy (saifoo.khan):

And don't even think about it if you're a doctor.

OpenStudy (geometry_hater):

Its logic so nothing in the universe will make 2 + 2 = 5

OpenStudy (anonymous):

If zero or one are defined oddly, this could be the case. And, why would you report this question? For example, this is possible in the ring \(\mathbb{Z}/0\mathbb{Z}\)

OpenStudy (anonymous):

A mental one.

OpenStudy (lgbasallote):

mathematicians believe in the impossible and improbable

OpenStudy (anonymous):

Since the only element in the ring is \(\{0\}\), and the other numbers are defined in terms of identities, every number could actually be a redundant symbol for such.

OpenStudy (anonymous):

well, this would happen if you assumed 4 by the symbol of 5 :)

OpenStudy (anonymous):

Correction to my post, I meant the ring \(\mathbb{Z}/1\mathbb{Z}\)

Parth (parthkohli):

This is possible on a planet where people don't know Mathematics.

OpenStudy (across):

If the symbol \(2\) represented the quantity \(2.5\), then this would be true.

OpenStudy (tyteen4a03):

I think the author is expecting a reference from 1984, so here's one: In Oceania, 2 + 2 = 5.

OpenStudy (nincompoop):

Orwell?

OpenStudy (maheshmeghwal9):

@nincompoop whose answer r u going to trust at? Sorry but i oppose such questions. There is no-one who can give correct answer.

OpenStudy (tyteen4a03):

@nincompoop Yes, Orwell's novel 1984.

OpenStudy (nincompoop):

2 people have demonstrated that it is possible and 1 person made a good reference. May I just add the terms "precision and accuracy?"

OpenStudy (tyteen4a03):

You will have to define the symbol 2, +, = and 5.

OpenStudy (anonymous):

Yes, it is possible. Consider \(\mathbb{Z}\) as the integers, and do not define that, in any manner, \(0\ne 1\). Therefore, \(\mathbb{Z}\) is indifferentiable from \(\{0\}\), so, any operation conducted under such ring given any number is valid, and equal, for all numbers in \(\mathbb{Z}\), given those axioms.

Parth (parthkohli):

If you go back to basics, then I'd say that the definitions would do so. If 2 was to be defined as 2.5, then yes.